Compound amount


use the compound interest formula to calculate the compound amount on 300,000 invested at 8% interest compounded quarterly for 5 years.


1
Expert's answer
2021-03-17T07:59:40-0400
"A=P(1+\dfrac{r}{n})^{n(t)}"

Given

"P=300000"

"r=0.08"

"n=4"

"t=5"


"A=300000(1+\dfrac{0.08}{4})^{4(5)}"

"A=445784.22"